What acceleration will result when a 12 N force to the right and 2.0N frictional force is applied to a 3.0 kg block of iron?
1
Expert's answer
2021-07-27T02:05:04-0400

The Newton's second law says

"a=\frac{F_{\rm net}}{m}=\frac{(12-2.0)\:\rm N}{3.0\:\rm kg}=3.3\:\rm m/s^2"